OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined: 7/4z+1

jimthompson5910 (jim_thompson5910):

You CANNOT divide by zero So the denominator 4z+1 CANNOT be equal to zero If it were, then 4z+1 = 0, which would mean that z = -1/4 So the value z = -1/4 makes the denominator zero and the expression undefined.